Two tuning forks having frequencies of 428 and 454 Hz are struck simultaneously. What average frequency will you hear, and what will the beat frequency be?

Average frequency:    
Beat frequency:   
Answer #1

average frequency heard = f1 + f2 / 2

= 428 + 454 / 2

= 441 Hz

beat frequency = f2 - f1

= 454 - 428 = 26 Hz